OpenStudy (anonymous):

Bring 5(3)^1/4 to an entire surd

OpenStudy (kropot72):

Use the following rule: \[a ^{\frac{m}{n}}=\sqrt[n]{a ^{m}}\]

OpenStudy (anonymous):

i think i have an answer

OpenStudy (anonymous):

if i rewrite 5 as 5^4/4 i have the (5^4*3)^1/4

OpenStudy (anonymous):

or (1875)^1/4

OpenStudy (kropot72):

\[5\times 3^{\frac{1}{4}}=5\times \sqrt[4]{3}\]